OpenStudy (anonymous):

A prism with a regular hexagonal base and a height of 3 in. has a volume of 288 square roots of 3 cu in. what is the length of each edge of the base?

OpenStudy (callisto):

Volume = Base area x height \[288\sqrt3 = base \ area \times 3\]\[96\sqrt3 = base \ area \]For a regular hexagon, it is made up of 6 equi. triangles. Height of each small triangle ----------------------- = tan 60 (x/2) Height of each small triangle = (x tan60) /2 Area of each smaller triangle = x [ (x tan60) /2 ] /2 = x^2 (sqrt3) /4 \[Area \ of \ hexagon = 6 \times [ x^2 (\sqrt3) /4 ] = 96 \sqrt3\]\[3x^2 (\sqrt3) = 192 \sqrt3\]\[x^2 = 64\] x =8 or x=-8 (rejected)